TMS Meeting Minutes, September 16, 2021

Attendees: Andrew, Breann, Heejoo, John, Matthieu, Olga, Trenton, Yang (Zoom Meeting)

Likely Misalignment between Fibers and Machine Channels
  • Issues with mismatched fibers to channels
  • Also, likely instabilities with the mirror cell magnetic mount plates
  • Leroy is searching for a fiber testing instrument to check for broken fibers
Recent Data Analysis from Heejoo
  • Heejoo prepared analysis based on the recent active TMS data from 20210915 and 20210916
  • Andrew contests that ch. 22 was not used, and so it should be removed from the analysis
  • Yang's code filters out channels using a sensitivity matrix
Active TMS during Co-Pointing
  • John thinks that there is a focus shift induced by adding in the U filter
  • However, it appears the theory was wrong
  • The filter focus offset of the U filter from DMS does NOT appear to be the cause of the issue
  • The issue appears to come up between each filter transition
  • Specifically, it's the biggest transition that causes the most problems
  • Andrew suggests potentially considering skipping beats on TMS while doing filter swaps, or simply not sending a correction when the filter is swapped
    • What if the TMS code were able to poll which filter was in LBC during each measurement
    • If the value of the filter is the same as for the previous measurement, measure
    • If the filter value changes, skip a measurement and pick up after stabilization
  • There doesn't appear to be a problem when it goes back down from the 0.4 filter shift, but it appears to crop up most during the largest focus transition
  • Also, it seems that the std. dev. filter is not working properly in this case but is misbehaving in these specific instances
    • The std. dev. filter should have seen hundreds of microns of change between measurements when the filters change and not collected measurements
  • When slewing to a new target and co-pointing, TMS threw in a bunch of garbage positions
  • Note, John changed the code to check for if the mirror is moving at the end of the measurement, but that doesn't appear to have fixed things
  • Andrew wonders if there is an issue with walleye causing a limitation in the allowable correction range
  • Being able to range balance the binocular is a necessary feature for TMS to be able to work peacefully with the telescope
  • However, John notes that TCS is already range balancing while in binocular mode
  • Regardless, TMS also needs to range balance to ensure we aren't pulling focus between the two sides
  • Despite having some issues with weird positioning, the TMS appears to have worked fairly well
Thermal Modeling
  • Andrew would like to see his thermal model implemented into the TMS code
  • John notes that this is best implemented as a Python script and that taking a consistent average of thermocouple data and finding the right reference could be challenging
  • Trenton works on the 3D modeling of the primary mirrors.

-- TrentonBrendel - 16 Sep 2021
Topic revision: r2 - 17 Sep 2021, HeejoChoi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback